# Forecasting from Clinical Textual Time Series: Adaptations of the Encoder and Decoder Language Model Families

### v2

@@ 4 Results
Table 1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1. Bold values indicate best in category amongst all models.
T2S2 (DeepSeek R1) sepsis-10 sepsis-100
F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
@@
DS-L3.3 70B 0.075 0.482 0.796 0.632 0.397 0.578 0.613 0.595
@@
ModernBERT-base 0.306 0.576 0.877 0.646 0.428 0.562 0.607 0.558
@@
ModernBERT-base 0.186 0.449 0.697 0.672 0.325 0.576 0.503 0.613
@@
ModernBERT-base 0.169 0.468 0.804 0.496 0.358 0.501 0.511 0.499
ModernBERT-large 0.133 0.349 0.534 0.494 0.271 0.509 0.277 0.502
Table 2: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from Llama 3.3-70B. Bold values indicate best in category amongst all models.
T2S2 (Llama 3.3) sepsis-10 sepsis-100
F1(1h) F1(1d) F1(1w) c-index F1(1d) c-index F1(1d) c-index
LLM-MLP head
DS-L3.3 70B 0.140 0.563 0.811 0.624 0.397 0.585 0.629 0.598
Prompting
L3.3 70B few-shot ordinal 0.064 0.460 0.729 0.313 0.480
Encoder-MLP head
ModernBERT-base 0.257 0.645 0.903 0.594 0.431 0.590 0.626 0.565
Encoder-masking-fine-tuned
ModernBERT-base 0.139 0.489 0.700 0.632 0.334 0.595 0.478 0.612
Encoder-masking-zeroshot
ModernBERT-base 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
4.1 Event forecast within next 24 hours: F1 performance
@@
Our results demonstrate a clear advantage of encoder-based models over LLMs in event forecasting, for both textual time-series annotations generated from DeepSeek-R1 (Table 1) and Llama-3.3-70B (Table 2) and particularly in short-term and medium-term prediction windows. The complete forecasting performance for all models have been shown in Table 6 and Table 7 in Appendix A. Encoder models, particularly those with a fine-tuned MLP head, achieve substantially higher F1 scores than the LLMs, reinforcing the effectiveness of encoder-based representations for forecasting.
@@
Among encoder models, different training strategies yield distinct performance trends. The MLP head fine-tuning approach consistently outperforms both fine-tuned masking and zero-shot masking models, particularly in long-horizonforecasting. ModernBERT-base achieves the highest overall F1 scores across all forecasting horizons, particularly at 168 hours, where it outperforms all other models. ModernBERT-large follows closely but does not consistently surpass its base variant, suggesting diminishing returns with increased model size. Meanwhile, zero-shot masking models, particularly standard BERT and RoBERTa, fail to make meaningful predictions, often yielding near-zero F1 scores at the 1-hour mark. However, DeBERTa-small and ModernBERT-base exhibit relatively better zero-shot performance, indicating that certain architectures generalize better without task-specific training.
Performance trends also vary across different datasets. Across all models, F1 scores improve as the forecasting window increases, with 1-hour predictions being the most challenging and 168-hour predictions yielding the highest scores. This suggests that longer-term trends are more predictable, whereas immediate event forecasting remains difficult, possibly due to the sparsity and volatility of short-term event occurrences. Comparing internal and external datasets, models perform worse on external validation sets (sepsis-10 and sepsis-100) than on the internal T2S2 test set, with a particularly noticeable drop in F1 at the 1-day horizon. This indicates potential generalization challenges, as models trained on internal data struggle to adapt to distributional shifts in clinical settings. Notably, ModernBERT models maintain a relative performance edge even in external datasets, reinforcing their robustness across different domains.
4.2 Temporal ordering of forecasted events: Concordance
In addition to event occurrence prediction, models are evaluated on their ability to correctly rank the order of upcoming events, measured through concordance (c-index) for both textual time-series annotations generated from DeepSeek-R1 (Table 1) and Llama-3.3-70B (Table 2). The complete forecasting performance for all models have been shown in Table 6 and Table 7 in Appendix A.
Similar to F1 trends, encoder-based models significantly outperform LLMs, with LLMs failing to achieve reliable ordinal rankings even when fine-tuned or prompted. Within encoder models, different training approaches show distinct performance characteristics: MLP-head fine-tuned models dominate in event occurrence classification (F1) but do not always yield the best concordance values. Instead, fine-tuned masking models, particularly ModernBERT-large, achieve the highest concordance scores, suggesting that this approach is better suited for ordering rather than binary classification.
Performance differences across datasets also highlight the challenges of generalization. While internal (T2S2) test results show relatively strong concordance for fine-tuned models, c-index values drop considerably on external datasets, particularly for models that are not ModernBERT-based. Among the best-performing models, ModernBERT-large (fine-tuned masking) achieves the highest c-index across both internal and external datasets, reinforcing its strength in capturing event order relationships. However, zero-shot masking models continue to struggle, mirroring their poor F1 performance. Additionally, while MLP-head fine-tuned models maintain high F1 scores, they exhibit slightly lower concordance than fine-tuned masking models, particularly at the 168-hour horizon. This suggests that different optimization strategies may be required depending on whether the task prioritizes strict classification accuracy or relative event ordering.
Overall, the results confirm that ModernBERT architectures are the strongest performers across both F1 and concordance evaluations, with fine-tuned MLP-head models excelling in binary prediction tasks and fine-tuned masking models achieving superior ranking performance. However, all models face challenges in generalizing to external datasets, emphasizing the need for domain adaptation techniques to improve real-world forecasting robustness.
4.3 Survival analysis
Model Performance across Timepoints and Cohorts.
Table 3 presents the time-dependent concordance index (concordance_td) for all models across three evaluation cohorts—t2s2, Sepsis10, and Sepsis100—at three observation windows (0h, 24h, and 168h). For each cohort and timepoint, we report performance from the best-fitting survival model (selected based on validation concordance).
Across cohorts, we observe that providing additional observation time (e.g., from 0h to 168h) does not consistently improve predictive performance. In the t2s2 cohort, several models achieve their peak concordance at the earliest or intermediate timepoints (e.g., DeepSeek-R1-Llama-70B peaks at 0h, while Llama-3.3-70B-Instruct peaks at 24h), with some degradation at 168h. In Sepsis100, certain models benefit from longer observation (e.g., DeepSeek-R1-Distill-Llama-8B improves steadily across timepoints), but others decline or plateau. The Sepsis10 cohort exhibits high and uniform performance across models, particularly after 24h, suggesting potential ceiling effects or limited evaluation variability in that subset.
Instruction-tuned LLMs such as Llama-3.3-70B-Instruct, DeepSeek-R1-Distill- Llama-70B, and Llama-3.1-8B-Instruct consistently outperform conventional transformer baselines like bert-base-uncased or roberta-base. This trend is particularly notable in the t2s2 and Sepsis100 cohorts, where LLMs demonstrate stronger early signal detection at 0h and 24h. These results suggest that large-scale pretraining and instruction tuning confer advantages in extracting actionable clinical representations, even with limited temporal context.
Taken together, these findings highlight that strong early performance does not require extended observation windows and that pretrained instruction-following models can offer substantial benefits for survival prediction in real-world, noisy clinical text scenarios.
@@ 4.3 Survival analysis
Table 3: Concordance index (concordance_td) for each model across three test sets (t2s2, Sepsis10, Sepsis100), evaluated at three observation time thresholds (0h, 24h, 168h).
4.4 Sensitivity analysis
To better understand the robustness and design implications of our forecasting models, we conduct two complementary sensitivity analyses focused on the temporal structure and completeness of input sequences. First, we compare two strategies for ordering clinical events within textual time-series (TTS) inputs: text-ordered, where events follow their extraction sequence from the original case report narratives, and time-ordered, where events are chronologically sorted. This comparison assesses the relative benefits of preserving narrative flow versus enforcing temporal alignment. Second, we evaluate the impact of input sparsity by randomly masking portions of the clinical history, simulating scenarios with incomplete documentation or missing data. Together, these analyses shed light on the importance of chronological ordering and data completeness for both classification accuracy and temporal event ranking.
4.4.1 Time-ordered vs text-ordered training
We evaluated the impact of training on text-ordered versus time-ordered textual time-series (TTS) annotations. In the text-ordered condition, events in the textual time-series order follow the order in which they were extracted from the case reports using LLMs (DeepSeek-R1 and Llama-3.3-70B), while in the time-ordered condition, they are sorted chronologically in time. This comparison aimed to assess the impact of chronological versus narrative ordering of events on forecasting performance
@@ 4.4.1 Time-ordered vs text-ordered training
As shown in Table 4, time-ordered training generally leads to higher concordance (C-index), particularly for DeepSeek-R1 annotatiobs across both internal (T2S2) and external test datasets (sepsis-10 and sepsis-100). However, F1 scores are less consistently affected by the ordering strategy. In several cases—especially with Llama-3.3-70B annotations on external datasets, text-ordered training results in equal or better F1 scores, suggesting that preserving the narrative sequence can support accurate classification even when it does not follow temporal order. These results highlight a trade-off between chronological alignment and narrative structure in optimizing forecasting performance.
@@
Table 4: Comparison of text-ordered and time-ordered training for Deep Seek R1 and Llama 3.3 annotations on F1 scores and C-index values.
Dataset Metric DeepSeek-R1 Llama 3.3-70B
Text-ordered Time-ordered Text-ordered Time-ordered
T2S2 Test F1 (1hr) 0.290 0.306 0.311 0.257
F1 (1day) 0.569 0.576 0.648 0.645
F1 (7days) 0.876 0.877 0.898 0.903
C-index 0.659 0.672 0.586 0.632
Sepsis10 F1 (1day) 0.361 0.428 0.423 0.382
C-index 0.598 0.576 0.526 0.595
Sepsis100 F1 (1day) 0.615 0.607 0.662 0.616
C-index 0.609 0.613 0.541 0.612
4.4.2 Impact of Randomly Masking History on F1 and Concordance
Further, we systematically evaluate how randomly omitting portions of clinical history impacts both forecasting accuracy and the model’s ability to correctly order clinical events. Towards this end, we simulate incomplete patient histories by randomly masking input time steps at varying rates. We refer to this as the timestep drop rate (TDR), ranging from 0% (no masking) to 90% (severely truncated history). We assess the effect of this masking on both classification performance (F1 score at 1 day) and ranking performance (concordance index) across three datasets—T2S2-DS, Sepsis-10, and Sepsis-100—using both DeepSeek-R1 (Table 5) and Llama-3.3-70B annotations (Table 8 in Appendix B).
F1 scores decline progressively with increasing timestep dropout rate (TDR), with sharp reductions beyond 60% (Table 5). At TDR = 0%, the F1 scores for T2S2-DS, Sepsis-10, and Sepsis-100 are 0.576, 0.428, and 0.607, respectively. By TDR = 90%, these values fall to 0.423, 0.200, and 0.510, with Sepsis10 showing the steepest decline. Between 30% and 60% dropout, performance remains relatively stable for T2S2-DS and Sepsis100, while Sepsis10 exhibits a more pronounced drop at TDR = 60%. Similar trends are observed with Llama-3.3 annotations (Table 8 in Appendix B).
In contrast, concordance (c-index) remains largely consistent across dropout levels. T2S2-DS (Table 5) reaches its highest concordance (0.678) at TDR = 60%, while Sepsis-10 and Sepsis-100 vary only marginally. This suggests that while masking temporal history substantially reduces classification accuracy, it has limited impact on event ranking performance. Comparable patterns are observed for the Llama-3.3 models (Table 8).
Overall, DeepSeek models show notable degradation in F1 scores when more than 30% of history is removed—particularly for datasets sensitive to longer-term dependencies—whereas concordance scores remain relatively stable (Table 5 and Table 8).
@@ 4.4.2 Impact of Randomly Masking History on F1 and Concordance
Table 5: Effect of Randomly Masking History (DeepSeek R1 annotations) on F1 and Concordance. TDR : Timestep drop rate
TDR F1 (1 day) Concordance (c-index)
T2S2-DS Sepsis-10 Sepsis-100 T2S2-DS Sepsis-10 Sepsis-100
0% 0.576 0.428 0.607 0.672 0.576 0.613
15% 0.572 0.415 0.604 0.667 0.564 0.616
30% 0.548 0.357 0.591 0.672 0.542 0.619
60% 0.542 0.310 0.588 0.678 0.578 0.616
90% 0.423 0.200 0.510 0.661 0.561 0.614
5 Discussion and Conclusion
Our findings highlight the superiority of encoder-based models over decoder-based LLMs for event forecasting, emphasizing the limitations of autoregressive models in structured prediction tasks. Despite prompting and fine-tuning, LLMs struggle with short- and medium-term forecasting, reinforcing the need for models optimized for sequence encoding rather than generative token prediction.
@@ 5 Discussion and Conclusion
Among encoder models, ModernBERT-base and ModernBERT-large consistently perform best, with MLP-head fine-tuning excelling in F1 scores and fine-tuned masking models achieving higher concordance (c-index), particularly in long-term forecasting. This suggests that classification and ranking tasks benefit from distinct optimization strategies, and a hybrid training approach may further enhance performance.
The gradual improvement of F1 scores over longer time horizons indicates that event patterns become more predictable over time, whereas short-term forecasting remains challenging due to higher variability. This limitation suggests that incorporating temporal modeling techniques, such as recurrent or convolutional structures, could refine short-term predictions while maintaining strong long-range forecasting.
Performance drops on external validation datasets (sepsis-10 and sepsis-100) highlight generalization challenges, with models struggling to adapt to shifts in event distributions. While ModernBERT models exhibit relative robustness, domain-specific fine-tuning or adaptive learning strategies are needed to bridge this gap.
Our survival analysis results complement the forecasting findings by revealing that strong predictive performance is often achievable with limited observation windows. Several models reach peak concordance at early (0h) or intermediate (24h) timepoints rather than at the longest window (168h), particularly in the T2S2 and Sepsis100 cohorts. Instruction-tuned LLMs such as Llama-3.3-70B-Instruct and DeepSeek-R1-Distill-Llama-70B outperform traditional transformer baselines in these survival settings, suggesting that instruction tuning and large-scale pretraining confer meaningful advantages for modeling time-to-event outcomes from clinical narratives in the textual time series setup.
Sensitivity analyses further highlight the importance of temporal structure and data completeness. When comparing time-ordered and text-ordered training regimes, we find that time-ordered training generally improves concordance, particularly for DeepSeek-R1 annotations, while text-ordering can yield competitive or better F1 scores on external datasets. This underscores a trade-off between temporal alignment and narrative coherence, with implications for designing learning objectives based on downstream task priorities. Moreover, robustness experiments simulating documentation sparsity reveal that F1 scores degrade significantly as more clinical history is masked, particularly beyond a timestep dropout rate of 60%. In contrast, concordance remains relatively stable under these conditions, indicating that event ranking is less sensitive to partial history than event classification.
Overall, our results demonstrate that encoder-based models, particularly ModernBERT, are well-suited for event forecasting but require enhancements in short-term prediction and cross-domain generalization. In contrast, for survival analysis tasks, we find that instruction-tuned LLMs have more predictive power. Future work should explore hybrid training objectives that balance classification and ranking performance, incorporate temporal modeling architectures, and investigate domain adaptation and pretraining strategies that explicitly account for data sparsity and real-world variability that we observe in our clinical textual time series.
Acknowledgments
This research was supported in part by the Division of Intramural Research (DIR) of the National Library of Medicine (NLM), National Institutes of Health. This work utilized the computational resources of the NIH HPC Biowulf cluster.
@@ Acknowledgments
Tables 6 and 7 present the forecasting performance of our models on textual time-series annotations generated by DeepSeek-R1 and Llama-3.3-70B, respectively. We report standard forecasting metrics across multiple evaluation settings to provide a comprehensive view of model performance.
Table 6: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1. Bold indicates best in that particular category amongst all models.
@@
Table 7: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from Llama-3.3-70B.

### v3

@@ Introduction
Table 1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. All models are trained/fine-tuned on time-ordered annotations from either DeepSeek-R1 or Llama 3.3-70B. Bold values indicate best in category within each column group (refer to Tables A1 and A2 in Appendix A for detailed performance statistics on all models). Performance statistics for all models are presented in A1 and A2, respectively. Abbreviations: F1(1d)-10/100: F1 (1 day) for sepsis-10 and sepsis-100 respectively. c10/100: c-index for sepsis-10 and sepsis-100 respectively.
T2S2 (DeepSeek-R1) T2S2 (Llama 3.3-70B)
F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100 F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100
@@
DS-L3.3 70B 0.075 0.482 0.796 0.632 0.397 0.578 0.613 0.595 0.140 0.563 0.811 0.624 0.397 0.585 0.629 0.598
OLMO 32B Instruct 0.000 0.411 0.688 0.621 0.273 0.561 0.432 0.593 0.190 0.315 0.764 0.604 0.238 0.551 0.315 0.596
RedPajama-INCITE 7B Instruct 0.000 0.352 0.651 0.618 0.282 0.572 0.471 0.594
MediPhi-PubMed 0.000 0.268 0.751 0.626 0.169 0.561 0.210 0.597 0.000 0.507 0.801 0.623 0.277 0.564 0.534 0.600
@@
ModernBERT-base 0.306 0.576 0.877 0.646 0.428 0.562 0.607 0.558 0.257 0.645 0.903 0.594 0.431 0.590 0.626 0.565
Bioclinical ModernBERT-base 0.264 0.559 0.879 0.677 0.395 0.598 0.635 0.610 0.290 0.653 0.902 0.650 0.449 0.618 0.662 0.627
@@
ModernBERT-base 0.186 0.449 0.697 0.672 0.325 0.576 0.503 0.613 0.139 0.489 0.700 0.632 0.334 0.595 0.478 0.612
Bioclinical ModernBERT-base 0.166 0.450 0.692 0.676 0.336 0.604 0.499 0.655 0.129 0.481 0.733 0.653 0.331 0.615 0.475 0.658
@@
ModernBERT-base 0.169 0.468 0.804 0.496 0.358 0.501 0.511 0.499 0.132 0.525 0.840 0.556 0.352 0.503 0.511 0.501
Bioclinical ModernBERT-base 0.246 0.458 0.614 0.498 0.301 0.509 0.457 0.498 0.197 0.460 0.533 0.499 0.299 0.520 0.465 0.499
Results
In this section, we present results for our three main evaluation settings: event forecast within a subsequent time window, temporal ordering of forecasted events, and survival analysis. We then provide sensitivity analyses on the forecasting tasks, examining the effects of temporal ordering and historical context availability.
Forecasting Tasks
Event forecast within next 24 hours: F1 performance.
@@ Results
Our results show that encoder-based models outperform decoder-based LLMs in event forecasting across both DeepSeek-R1 and Llama-3.3-70B annotations (Table 1). Forecasting performance across all models is shown in Tables A1 and A2 (Appendix A). Encoder models, especially those with a fine-tuned MLP head, achieve substantially higher F1 scores than decoder LLMs, reinforcing the effectiveness of encoder-based representations for forecasting.
Among decoder models, we observe varied performance patterns. RedPajama-INCITE-7B-Instruct, with documented training that excludes PubMed, achieves 0.352 F1 at 24h—better than some decoder models like MediPhi-PubMed (0.268) and within the range of decoder performance, though below top performers like DeepSeek-Llama-3.3-70B (0.482). OLMO-32B-Instruct performs competitively (0.411), suggesting that open models without biomedical pretraining can still achieve reasonable results for temporal forecasting tasks.
Among encoder models, both general-purpose and biomedically-pretrained variants show strong performance. While ModernBERT-base achieves slightly higher F1 scores on internal test sets (0.576 vs 0.559), BioClinical-ModernBERT-base demonstrates superior concordance (0.677 vs 0.646 for DeepSeek-R1) and notably better generalization to external datasets, with F1 scores of 0.635 and 0.662 on sepsis-100 compared to 0.607 and 0.626 for the general-purpose variant. This suggests that while both ModernBERT architectures are effective, biomedical pretraining particularly enhances temporal reasoning and cross-dataset generalization.
@@
Different training strategies yield distinct performance patterns. The MLP head fine-tuning approach consistently outperforms both fine-tuned masking and zero-shot masking models, particularly in long-horizon forecasting. While ModernBERT-base achieves strong F1 scores, BioClinical-ModernBERT-base demonstrates the best overall performance when considering both accuracy and generalization. Zero-shot masking models, particularly standard BERT and RoBERTa, fail to make meaningful predictions at the 1-hour mark. However, ModernBERT variants exhibit relatively better zero-shot performance, with BioClinical-ModernBERT-base achieving notably higher scores in zero-shot settings (0.246 F1 at 1h), indicating that biomedical pretraining improves out-of-the-box temporal reasoning.
Performance trends remain consistent across forecasting windows. F1 scores improve as the window increases, with 1-hour predictions being most challenging and 168-hour predictions yielding highest scores. Models perform worse on external validation sets than internal T2S2 test sets, though BioClinical-ModernBERT models show the smallest performance degradation, maintaining their advantage in cross-dataset generalization.
Temporal ordering of forecasted events: concordance.
Encoder models consistently outperform decoders in correctly ranking the order of upcoming events, as measured by concordance (c-index; Table 1; complete results in Tables A1 and A2 for DeepSeek-R1 and Llama-3.3-70B annotations, respectively, with full details in Appendix A. Decoder models with potential biomedical exposure achieve modest concordance (0.618-0.632), while RedPajama-INCITE-7B-Instruct without PubMed data in its training corpus performs similarly, suggesting domain-specific knowledge is less critical here than for event prediction. The comparable performance of RedPajama-INCITE-7B-Instruct to other general models (with access to PubMed in their pretraining) also provides evidence that the T2S2 prediction task is distinct from raw PubMed text and assuages causal leakage concerns.
Biomedical pretraining in encoder models also demonstrates clear advantages, with BioClinical-ModernBERT-base attaining the highest concordance (0.677 with MLP head, 0.676 fine-tuned masking) and robust performance across datasets. MLP-head training yields superior F1 scores, while fine-tuned masking excels in concordance. BioClinical-ModernBERT-base maintains strong generalization on external datasets (c-index >0.60), suggesting that biomedical pretraining provides stability in temporal reasoning across dataset shifts.
Survival analysis: model performance across timepoints and cohorts.
Table 2 presents time-dependent concordance across evaluation cohorts (T2S2, sepsis-10, sepsis-100) at multiple observation windows (0h, 24h, 168h), selecting models based on validation concordance. General models, such as RedPajama-INCITE-7B-Instruct, show strong survival prediction capabilities even without biomedical exposure (0.76 concordance at 168h, sepsis-100). Meanwhile, biomedical encoders like BioClinical-ModernBERT-base consistently perform well, particularly early (0h).
Predictive performance does not universally improve with more observation time. T2S2 cohort models often peak at earlier windows, while sepsis-10 experiences ceiling effects. Instruction-tuned decoders excel overall, yet biomedical encoders remain competitive, particularly at admission (0h). Survival prediction thus benefits from both larger decoder models and domain-specific pretraining, each excelling in different temporal contexts.
Table 2: Survival analyses evaluated using time-dependent concordance index for each model across DeepSeek-R1 test set annotations (T2S2, Sepsis-10, Sepsis-100), evaluated at three observation time thresholds (0h, 24h, and 168h).
@@
Bioclinical ModernBERT-base 0.57 0.53 0.53 1.00 0.75 0.75 0.63 0.59 0.56
Bioclinical ModernBERT-large 0.54 0.55 0.54 0.75 0.75 0.75 0.61 0.57 0.49
@@
OLMO 32B Instruct 0.60 0.62 0.59 0.75 0.75 0.75 0.66 0.58 0.56
RedPajama-INCITE 7B Instruct 0.56 0.60 0.61 0.50 1.00 1.00 0.61 0.53 0.76
MediPhi-PubMed 0.54 0.62 0.55 1.00 1.00 1.00 0.62 0.58 0.61
Sensitivity analysis
Time-Ordered versus Text-Ordered Training.
@@
Table 3 shows time-ordering generally improves concordance, particularly for DeepSeek-R1 annotations, while F1 scores show mixed results. Text-ordering sometimes yields better external dataset performance with Llama-3.3-70B, suggesting that preserving the narrative sequence can support accurate classification even when it does not follow temporal order. These results highlight a trade-off between temporal alignment and narrative coherence for different tasks.
@@
Table 3: Comparison of text-ordered and time-ordered training for DeepSeek-R1 and Llama 3.3-70B annotations on F1 scores and c-index values.
F1(1h) F1(1d) F1(1w) c-index F1(1d)-10 c10 F1(1d)-100 c100
T2S2 (DeepSeek-R1)
Text-ordered 0.290 0.569 0.876 0.659 0.361 0.598 0.615 0.609
Time-ordered 0.306 0.576 0.877 0.672 0.428 0.576 0.607 0.613
T2S2 (Llama 3.3-70B)
Text-ordered 0.311 0.648 0.898 0.586 0.423 0.526 0.662 0.541
Time-ordered 0.257 0.645 0.903 0.632 0.382 0.595 0.616 0.612
Impact of timestep dropout on F1 and concordance.
Tables 4 and A3 (Appendix B) report the effect of timestep dropout on classification and ranking performance. F1 scores decrease monotonically with increasing TDR, with relatively mild degradation up to 30–60% dropout and sharper declines thereafter. For example, in the DeepSeek-R1 annotations, T2S2/sepsis-10/sepsis-100 drop from 0.576/0.428/0.607 at TDR = 0% to 0.423/0.200/0.510 at TDR = 90%, with sepsis-10 showing the largest relative decline. In contrast, concordance (c-index) remains stable across dropout levels (0.661–0.678), indicating that ranking-based evaluations are less sensitive to history truncation. These trends are consistent for Llama-3.3 annotations (Appendix B), suggesting that ranking objectives capture temporal relationships that are more robust to missing historical information than binary classification accuracy of predicting whether an events happens within a time window.
@@
Table 4: Effect of Randomly Masking History (DeepSeek-R1 annotations) on F1 and Concordance.
Timestep drop rate F1(1d) c-index F1(1d)-10 c10 F1(1d)-100 c100
0% 0.576 0.672 0.428 0.576 0.607 0.613
15% 0.572 0.667 0.415 0.564 0.604 0.616
30% 0.548 0.672 0.357 0.542 0.591 0.619
60% 0.542 0.678 0.310 0.578 0.588 0.616
90% 0.423 0.661 0.200 0.561 0.510 0.614
Discussion
@@
Our findings highlight the superiority of encoder-based models over decoder-based LLMs for event forecasting, emphasizing the limitations of autoregressive models in structured prediction tasks. Among encoder models, BioClinical-ModernBERT-base consistently performs best, achieving superior concordance (0.677 vs 0.646) and better external generalization despite slightly lower internal F1 scores than general-purpose ModernBERT-base. MLP-head fine-tuning excels in F1 scores while fine-tuned masking models achieve higher concordance, particularly in long-term forecasting. This suggests that classification and ranking tasks benefit from distinct optimization strategies.
The consistent advantage of biomedically-pretrained models across tasks deserves special attention. While general-purpose models achieve strong internal performance, biomedical pretraining provides crucial benefits for real-world deployment: BioClinical-ModernBERT variants show the smallest performance degradation on external datasets and maintain higher zero-shot capabilities (0.246 F1 at 1h vs near-zero for standard BERT/RoBERTa). This suggests that domain-specific pretraining enhances not just accuracy but also robustness and generalization—critical factors when considering clinical deployment where distribution shifts are common.
The gradual improvement of F1 scores over longer time horizons indicates that event patterns become more predictable over time, whereas short-term forecasting remains challenging due to higher variability. Performance drops on external validation datasets highlight generalization challenges, though ModernBERT models exhibit relative robustness. For survival analysis, instruction-tuned LLMs like Llama-3.3-70B-Instruct outperform traditional transformer baselines, with several models achieving peak concordance at early timepoints rather than with extended observation.
Our sensitivity analysis reveals several trade-offs: time-ordered training generally improves concordance while text-ordering can yield better F1 scores on externally-annotated datasets. Robustness experiments show F1 scores degrade significantly beyond 60% timestep dropout, while concordance remains stable, indicating event ranking is less sensitive to partial history than event classification.
Methodological Contributions and Potential Impact. Our framework demonstrates how narrative clinical texts can be systematically converted into structured temporal representations for forecasting tasks. The ability to extract temporally structured insights from unstructured clinical text could potentially support clinical decision-making, particularly in settings with limited access to structured data or specialized expertise.
Importantly, while consumers frequently consult LLMs about health risk via prompting in chats, this study demonstrates that the prompting approach performs substantially worse in risk prediction than alternative approaches, at least with respect to precision/recall/F1, with prompted LLMs achieving at best 0.460 F1 at 24h compared to 0.653 for fine-tuned encoders. Our study highlights several alternative approaches and characterizes their relative performance strengths and weaknesses. In addition, our finding that concordance remains stable with 60% missing context suggests robustness to incomplete documentation scenarios, which highlights the degradation pattern in performance as it relates to context availability.
Our approach also demonstrates that our language model systems can reliably capture the temporal reasoning that clinicians use in practice based on their clinical documentation. However, additional work would be required to validate performance on real-time clinical data and demonstrate measurable impact on patient outcomes in deployment. Further discussion on societal impact of our work is in Appendix I.
Limitations and Future Directions. Our study has important limitations to consider. First, the pipeline relies on case reports from the PubMed Open Access (PMOA) corpus, which often reflect rare or atypical presentations and differ from routine clinical notes (e.g., progress notes, discharge summaries), potentially limiting the model’s generalizability to real-world health system texts. Compared to routine clinical notes, case reports offer rich, temporally structured narratives with explicit clinical rationale, making them especially useful for evaluating model performance under sparse documentation. This study develops a general framework for textual time series forecasting, using case reports as an interpretable, temporally rich foundation for method development with potential extensions to real-world clinical corpora like MIMIC discharge summaries.
Second, while we focused on sepsis due to its clinical relevance and prevalence in case reports, our framework is fundamentally disease-agnostic. Preliminary results from over 125K PMOA case reports across various diagnostic conditions show promising generalization, with future work planned on broader diagnostic categories and real-world corpora such as MIMIC-IV discharge summaries.
Third, despite efforts to mitigate causal leakage during training and evaluation (e.g., timestamp extraction, temporal masking, and use of external test sets), pretrained language models may have been exposed to PMOA content. Although these models were not trained on our derived (event, time) sequences or temporally reordered data, it is possible that the original case report narratives were included in their pretraining data. Our pipeline applies a series of transformations—including temporal reordering and masking of future events beyond a prediction cutoff—which fundamentally alter the input representation and enforce a causal structure that is not present in the original documents. This creates a distinct prediction setup that requires models to reason over partial, temporally grounded inputs rather than complete narratives. To further assess the impact of pretraining exposure, we include results from RedPajama-INCITE-7B-Instruct, an open-source model with a fully documented, PubMed-free training corpus. Its comparable performance to other decoder models (0.352 F1 at 24h) and similar concordance (0.618) to models potentially exposed to PubMed provides evidence that our T2S2 prediction task is sufficiently distinct from raw PubMed text, helping to mitigate causal leakage concerns. A detailed discussion on the scope of using open-source models for avoiding potential data leakage has been presented in Appendix H.
Finally, our analysis is limited by the modest size of the high-cost manual annotations collected as a gold standard to calibrate and validate our LLM extraction pipeline. Expanding gold-standard validation across diverse diagnoses could assess the generalizability of our approach and aid in evaluating model robustness, and is a direction for future work.
Supplementary Material
This supplementary material provides detailed information organized as follows:
  • Appendix A: Complete forecasting performance tables for all models on DeepSeek-R1 and Llama-3.3-70B annotations
  • Appendix B: Sensitivity analysis of masking history on Llama-3.3-70B annotations
  • Appendix C: Technical details of encoder-based forecasting methods, including MLP-head and masking approaches
  • Appendix D: Survival modeling framework using language model embeddings and the hyperparameters used
  • Appendix E: Computing infrastructure used for experiments, including hardware specifications, memory requirements, operating system, and software environment
  • Appendix E: LLM prompt template for extracting clinical time series from case reports
  • Appendix E: Prompting strategies for zero-shot and few-shot forecasting
  • Appendix H: Analysis of open-source models to address data leakage concerns
  • Appendix I: Broader societal impact and ethical considerations
@@
Table A1: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. Models italicized represent the open-source and medically fine-tuned models. All models are trained/fine-tuned on time-ordered annotations from DeepSeek-R1. Bold indicates best in that particular category amongst all models.
@@
OLMO 32B Instruct 0.000 0.411 0.688 0.621 0.273 0.561 0.432 0.593
RedPajama-INCITE 7B Instruct 0.000 0.352 0.651 0.618 0.282 0.572 0.471 0.594
MediPhi-PubMed 0.000 0.268 0.751 0.626 0.169 0.561 0.210 0.597
@@
OLMO 32B 0-shot 0.018 0.333 0.652 0.322 0.430
OLMO 32B 0-shot ordinal 0.006 0.323 0.685 0.325 0.443
OLMO 32B few-shot 0.014 0.316 0.630 0.219 0.423
OLMO 32B few-shot ordinal 0.013 0.315 0.651 0.247 0.422
@@
Bioclinical ModernBERT-base 0.264 0.559 0.879 0.677 0.395 0.598 0.635 0.610
Bioclinical ModernBERT-large 0.313 0.562 0.868 0.671 0.399 0.491 0.618 0.607
@@
Bioclinical ModernBERT-base 0.166 0.450 0.692 0.676 0.336 0.604 0.499 0.655
Bioclinical ModernBERT-large 0.165 0.422 0.696 0.652 0.325 0.626 0.425 0.603
@@
Bioclinical ModernBERT-base 0.246 0.458 0.614 0.498 0.301 0.509 0.457 0.498
Bioclinical ModernBERT-large 0.244 0.446 0.596 0.479 0.364 0.502 0.416 0.497
Table A2: Forecasting performance (event occurrence: F1 and correct event ordering: concordance-index) of the ensuing $k=8$ events. Models italicized represent the open-source and medicallty fine-tuned models. All models are trained/fine-tuned on time-ordered annotations from Llama-3.3-70B.
@@
OLMO 32B Instruct 0.190 0.315 0.764 0.604 0.238 0.551 0.315 0.596
RedPajama-INCITE 7B Instruct 0.127 0.459 0.705 0.620 0.257 0.586 0.491 0.594
MediPhi-PubMed 0.000 0.507 0.801 0.623 0.277 0.564 0.534 0.600